OpenStudy (anonymous):

An article is bought for $7000 and sold 5 years later for $650. Find the depreciation formula, and the value of the article after 2 years. Let y be the value in dollars, and t be time in years.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: this is a linear equation and it goes through the two points (0,7000) and (5,650)

OpenStudy (anonymous):

wait so thats the answer?

OpenStudy (anonymous):

its asking find the depreciation formula

OpenStudy (anonymous):

and wants me to write it like this y= _t + _

OpenStudy (anonymous):

and then the value after 2 years =?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you have to find the equation first

jimthompson5910 (jim_thompson5910):

do you know how to find the equation through 2 points?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

is it y= mx+b

jimthompson5910 (jim_thompson5910):

what's the slope of the line?

OpenStudy (anonymous):

not sure

jimthompson5910 (jim_thompson5910):

use the slope formula m = (y2 - y1)/(x2-x1)

OpenStudy (anonymous):

yea but how this isnt a graph or it does not give us coordinates

jimthompson5910 (jim_thompson5910):

the two points are (0,7000) and (5,650)

OpenStudy (anonymous):

which one is y2

OpenStudy (anonymous):

is 0 x1

jimthompson5910 (jim_thompson5910):

y2 is the second y coordinate y2 = 650

OpenStudy (anonymous):

whats 5 y1

OpenStudy (anonymous):

(6d0-0)

OpenStudy (anonymous):

650

OpenStudy (anonymous):

wwhats the answer

jimthompson5910 (jim_thompson5910):

y1 is 7000

jimthompson5910 (jim_thompson5910):

so y2 - y1 = 650 - 7000 = ???

OpenStudy (anonymous):

-6350

jimthompson5910 (jim_thompson5910):

x2 - x1 = ???

OpenStudy (anonymous):

5-0

jimthompson5910 (jim_thompson5910):

which equals what

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

Now divide the two differences: -6350/5 = -1270

jimthompson5910 (jim_thompson5910):

The slope is -1270

jimthompson5910 (jim_thompson5910):

m = -1270

OpenStudy (anonymous):

it doesent ask for the slope though

jimthompson5910 (jim_thompson5910):

y = mx + b y = -1270x + b ... plug in m = -1270 7000 = -1270*0 + b ... plug in (x,y) = (0,7000) solve for b

OpenStudy (anonymous):

7000=-1270*0+what

OpenStudy (anonymous):

-1270 is this t?

jimthompson5910 (jim_thompson5910):

no that's the slope

OpenStudy (anonymous):

because its asking y=_t+_

jimthompson5910 (jim_thompson5910):

yeah I know

jimthompson5910 (jim_thompson5910):

7000 = -1270*0 + b solve for b

OpenStudy (anonymous):

how

OpenStudy (anonymous):

is it 7000/-1270

jimthompson5910 (jim_thompson5910):

-1270*0 = ??

OpenStudy (anonymous):

o

jimthompson5910 (jim_thompson5910):

7000 = -1270*0 + b 7000 = 0 + b b = ??

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

-1270?

jimthompson5910 (jim_thompson5910):

0 + b turns into b since adding 0 to some number doesn't change it

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

ok so whats b

jimthompson5910 (jim_thompson5910):

7000 = 0 + b you tell me

OpenStudy (anonymous):

7000

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

m = -1270 b = 7000

jimthompson5910 (jim_thompson5910):

y = mx + b y = -1270x + 7000 y = -1270t + 7000

OpenStudy (anonymous):

so then we solve to get after 2 years?

jimthompson5910 (jim_thompson5910):

plug in t = 2 and evaluate

OpenStudy (anonymous):

so -1270 times 2 +7000= price after 2 years

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

4,460?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thanks for explaining it actually now im kinda getting it :)

jimthompson5910 (jim_thompson5910):

you're welcome